Dr. Al Kali, an environmental chemist, measures the pH of numerous samples of water collected from a polluted lake. His data are normally distributed with a mean pH of 8.25 and a standard deviation of 0.35. What should Dr. Kali report as the margin of error for the 90% confidence interval for his data?
1
Expert's answer
2011-05-19T07:15:16-0400
Let X ne the pH of numerous samples of water collected from a polluted lake.
By assumption X has normal distribution N(8.25,0.35).
Then Z=((X-8.25)/0.35) has normal distribution N(0,1).

Then margin of error is equal to
δ = 0.35 * z, where z is such that
P(-z< Z < z) = 0.9
If Φ is a c.d.f. for N(0,1), then
z= - Φ-1 ((1-0.9)/2)= -Φ-1 (0.05)= 1.645

Hence
δ = 0.35 * 1.645 = 0.576